OpenStudy (anonymous):

Can someone please help me with an Algebra assignment? Gordon is evaluating some stocks he wants to purchase. You can see the function of the stock price below. The variable x represents days.

OpenStudy (anonymous):

This is the table that came with the assignment.

OpenStudy (kc_kennylau):

And what's the question?

OpenStudy (anonymous):

Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements.

OpenStudy (anonymous):

there are a couple others, I'm just not sure where to start.

OpenStudy (anonymous):

No, it says D.

OpenStudy (anonymous):

I think so

OpenStudy (tylerd):

the lowest starting price is 22

OpenStudy (tylerd):

so the second lowest would have to be between 22 and 25

OpenStudy (anonymous):

okay, how to i put that into my function. where does it go?

OpenStudy (tylerd):

the function is in the form of an exponential function f(x)=ab^x so choose a number between 22 and 25, say 23 f(x)=23(b)^x

OpenStudy (nincompoop):

you need to know compounding interest

OpenStudy (tylerd):

im not sure if they do need to know, looks like regular exponential functions to me.

OpenStudy (nincompoop):

whenever the number of days is not indicated, we use the e solution

OpenStudy (anonymous):

okay, so what is the e solution then?

OpenStudy (nincompoop):

that is something you need to figure out

OpenStudy (nincompoop):

read up on compounded and continuous interests

OpenStudy (nincompoop):

do you need a textbook or something?

OpenStudy (anonymous):

i can google it

OpenStudy (tylerd):

did your teacher go over any compound interest?

OpenStudy (anonymous):

Technically, it's not my teacher I do school online and i don't remember reading that in the chapter.

OpenStudy (tylerd):

ya

OpenStudy (tylerd):

i dont think thats what they are asking.

OpenStudy (tylerd):

but you should learn it anyway

OpenStudy (anonymous):

okay

OpenStudy (tylerd):

but for this problem i think something like \[f(x)=23(1.06)^x\] would probably work

OpenStudy (anonymous):

okay, there are actually two more questions, but for me this subject is kind of confusing. You think you can help?

OpenStudy (anonymous):

And thanks so much, I can kinda see how that function would work

OpenStudy (anonymous):

I had to read this a couple of times. A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x).

OpenStudy (kc_kennylau):

Notice that g(x) is a constant.

OpenStudy (kc_kennylau):

f(x)+g(x) = 23*1.06^x + (-6)

OpenStudy (kc_kennylau):